# Scalable Graph Attention‐Based Instance Selection via Mini‐Batch Sampling and Hierarchical Hashing

### v2

@@ 1 Introduction
Traditional IS methods rely mostly on nearest neighbor computations, falling into three groups: condensation (keeping instances near decision boundaries), edition (removing noisy instances), and hybrid approaches (Cano et al., 2003; Garcia et al., 2012; Saha et al., 2022). While these work well for moderately sized datasets, they struggle to capture complex relationships in high-dimensional spaces. Their need for pairwise distance calculations results in quadratic complexity that becomes prohibitive for large-scale datasets, often taking several hours of preprocessing time on datasets with millions of instances.
Graph-based methods offer a promising alternative by representing data relationships through node connections (Goyal and Ferrara, 2018). These approaches can effectively capture structural information through proximity graphs and neighborhood networks, revealing dataset patterns that distance-based methods might miss. Recent advances in graph sampling and reduction (Chen et al., 2022a) show their potential to handle large-scale computational challenges while retaining key data characteristics.
@@
Graph Attention Networks (GATs) (Veličković et al., 2017) present an unexplored direction for IS through their ability to learn importance weights for node neighborhoods dynamically. The attention mechanisms in GATs can compute instance importance based on local structural properties, providing an interpretable approach to identify informative data points. However, two questions remain: (1) How can we efficiently construct meaningful graph representations from large tabular datasets? (2) How can we utilize attention mechanisms to identify truly informative instances while keeping computational costs feasible?
@@ 2.1 Instance Selection
Traditional IS methods primarily rely on nearest neighbor computations to evaluate instance relationships and importance, falling into three main categories: condensation (retaining instances near decision boundaries), edition (removing noisy or borderline instances), and hybrid approaches (Garcia et al., 2012). While various improvements have been proposed, including density-based and scalability-focused approaches (Carbonera and Abel, 2018; García-Osorio et al., 2010), recent work has introduced several innovative methodologies that enhance or reformulate how instance relationships are evaluated.
Density-based approaches have evolved from local to global analysis, as demonstrated by GDIS and EGDIS (Malhat et al., 2020), which evaluate instance importance through global density patterns and neighbor relationships. The ranking-based RIS algorithm (Cavalcanti and Soares, 2020) introduces a scoring mechanism that considers relationships between instances, using normalized exponential transform of distances weighted by class membership. More theoretically grounded approaches such as BDIS (Chen et al., 2022b) employ Bayesian decision theory to categorize instances as reducible, irreducible, or deleterious, while leveraging percolation theory to identify local homogeneous clusters—groups of instances with the same label connected by nearest neighbors, where the cluster size reflects the probability of class consistency. Alternative frameworks like CIS (Moran et al., 2022) employ reinforcement learning and clustering, where an agent learns to select valuable instance clusters through Q-learning and curiosity-driven exploration. Recent work has explored novel theoretical frameworks, including approval-based multi-winner voting (Sánchez-Fernández et al., 2024), which models instances as both voters and candidates in an election system, and evidence theory-based approaches like EIS (Gong et al., 2021), which evaluates instance importance through mass functions and Dempster’s rule of combination.
@@ 2.2 Graph-based Selection
Graph reduction techniques (Chen et al., 2022a), while successful in general graph summarization tasks, have seen limited application in IS. As shown in Figure 2, these methods merge nodes or modify edge structures. When applied to IS, these approaches typically require mapping reduced nodes back to original instances, often by selecting representatives from merged clusters. This indirect correspondence makes these methods less suitable for IS tasks compared to graph sampling approaches that maintain direct relationships to the original instances.
@@
The effectiveness of graph-based methods for IS has been demonstrated by Rustamov et al. (2024b), who evaluated multiple graph-based approaches across diverse classification datasets, showing their potential to achieve substantial data reduction while maintaining model performance. In our previous work (Rustamov et al., 2024a), we explored using GATs for IS. While this approach showed promise in identifying important instances through attention mechanisms, it was limited by its reliance on confidence scores as importance measures and faced significant scalability issues with large datasets due to its inefficient chunking-based graph construction that still required substantial memory overhead. In this paper, we address these limitations by introducing efficient graph construction methods—distance-based mini-batching and LSH approaches—that significantly reduce computational complexity, alongside a dual-perspective importance scoring mechanism that leverages attention weights to more effectively identify informative instances.
@@ 2.3 Scaling Challenges in Graph-based Methods
While graph-based methods show promise for IS, they face significant scaling challenges as datasets grow in size and dimensionality. The primary bottleneck lies in the graph construction phase, particularly for k-nearest neighbor (KNN) graphs that form the foundation of many graph-based approaches. Constructing KNN graphs requires computing pairwise distances between instances, resulting in $O(n^{2})$ computational complexity and substantial memory requirements for storing the resulting adjacency matrices (Li et al., 2024).
@@ 3.1.1 Distance-based Mini-Batching Method
[…]
@@ 3.1.2 Locality-Sensitive Hashing-based Method
Figure 5: High-level overview of the Multi-Level LSH (ML-LSH) approach. ML-LSH processes instances through multiple hierarchical levels (Level $1$ to $M$), where each level follows a four-step process: (1) Random projections using hash functions, (2) Hash code generation (e.g., $h(x)=\text{sign}(P^{T}x)$), (3) Bucket assignment grouping similar instances, and (4) Edge connection within buckets. Lower levels (Level $1$) capture coarse relationships using fewer hash tables and projections, while higher levels (Level $M$) employ exponentially increasing hashing parameters to refine local similarities. The final graph combines edge sets from all levels, with each level treated independently to capture different relationship granularities.
@@
[…]
@@
[…]
@@ 3.2 Graph Attention Network (GAT)
GATs provide an ideal foundation for IS due to their ability to learn adaptive node relationships through attention mechanisms. While traditional GATs excel at capturing local structural patterns, IS requires a more nuanced understanding of both local and global data relationships. We build upon the foundational GAT architecture (Veličković et al., 2017) by introducing four key enhancements: (1) head-wise attention amplification through learnable scaling factors, (2) diversity-promoting attention that encourages distinct patterns across attention heads, (3) multi-level/multi-view attention mechanisms that process our graph representations at multiple scales, and (4) cross-level attention via Nyström approximation for efficient global information exchange.
@@ 3.3.2 Selection Strategies
Table 1: Dataset statistics with domain, number of instances ($N$), features ($D$), and classes ($K$). Dataset size indicators: small ($N<10,000$), medium ($00010,000\leq N\leq 100,$ 000), large ($N>100,000$).
@@ 4.1 Datasets and Models
This study employs diverse classification datasets from the UCI Machine Learning Repository (Kelly et al., 2023) to evaluate our proposed GAIS method. Table 1 summarizes these datasets, including their domain, number of instances ($N$), features ($D$), and classes ($K$). The datasets vary widely in size (from 303 to five million instances), dimensionality (two to 784 features), and class distributions (balanced and imbalanced), spanning domains such as healthcare, game theory, and bioinformatics. We evaluate our method using Random Forest and Logistic Regression as representative non-linear and linear classifiers, respectively.
@@ 4.5 Evaluation Metrics
Accuracy $AC$ = (TP + TN) / (TP + TN + FP + FN)
Precision $PR$ = TP / (TP + FP)
Recall $RE$ = TP / (TP + FN)
@@ 5 Results and Discussions
Method $AC$ $PR$ $RE$ $F1$ $E$ $E_{F1}$ $R$
Small Datasets ($<10,000$)
@@ 5.1 Comparing GAIS Variants
For large datasets, the DM method demonstrates the highest efficiency with the lowest construction and training times (GCT = 2.55s, GTT = 10.31s), making it particularly suitable for large-scale applications where computational resources are limited. The more complex MVML-LSH and CMVML-LSH methods show significantly higher computational requirements (GCT $>$ 35s, GTT $>$ 80s), though this should be weighed against their enhanced performance metrics. Importantly, the original ML training time (70.64s) is substantially higher than the training time on selected instances (0.28-0.82s), highlighting the significant computational savings achieved through IS, regardless of the chosen method.
@@ 5.3 Component Analysis
Figure 8 illustrates the relationship between mini-batch configuration and test effectiveness in the DM method for large datasets. We test configurations up to $w_{\max}$ instances per mini-batch to identify natural performance plateaus. The results demonstrate that both the number of mini-batches and instances per batch have a positive impact on effectiveness, though this improvement shows diminishing returns. As the number of mini-batches increases from 1 to 9, the effectiveness consistently improves across all instance sizes, with the most substantial gains observed in the transition from 1 to 5 mini-batches. Similarly, larger batch sizes generally yield better effectiveness, though the improvement plateaus around $7000$ instances, particularly with higher numbers of mini-batches. This plateau effect is most noticeable when using 7-9 mini-batches, suggesting that beyond these configurations, additional computational resources might not yield proportional performance benefits. These findings provide practical guidance for implementing the DM method, indicating that moderate configurations (5-7 mini-batches with $5000-7000$ instances per batch) might offer an optimal balance between computational efficiency and effectiveness. The empirical analysis demonstrates that effectiveness plateaus around $7000$ instances per mini-batch, providing direct guidance for $w_{\max}$ selection. This plateau effect validates that computational bounds can be set without sacrificing performance.
Figure 8: Impact of mini-batch configuration on test effectiveness in the distance-based mini-batch (DM) method for large datasets ($>$100K instances). Results show diminishing returns beyond 7,000 instances per mini-batch and 5-7 mini-batches, providing empirical guidance for setting the computational constraint $w_{\max}$ while maintaining performance.
@@
Figure 9: Effectiveness versus sampling percentage by dataset size categories: small ($<$10K instances), medium (10K–100K instances), and large ($>$100K instances). Performance plateaus occur at 5%, 10%, and 40% sampling for large, medium, and small datasets respectively, with the ”Diminishing Returns Zone” indicating where additional sampling provides minimal benefit.
@@
Method Sep $\uparrow$ Pur $\uparrow$ $R@5$uparrow $\rho$uparrow $\bar{B}$downarrow
Small Datasets ($<10,000$)
@@
Figure 11: Multi-level and multi-view analysis showing relative improvement (in %) in test effectiveness for LSH-based graph construction variants. (a) Multi-level analysis: Improvement from using hierarchical levels compared to single-level baseline, across small ($<$10K instances), medium (10K–100K instances), and large ($>$100K instances) datasets. (b) Multi-view analysis: Improvement from using multiple views compared to single-view baseline. Solid lines show average improvements across all datasets in each size category, while dashed lines represent improvements for datasets performing below median effectiveness within each category to highlight challenging datasets.
[…]
@@
Table 6: Ablation study of GAT enhancements for distance-based mini-batching (DM) and multi-level locality-sensitive hashing (ML-LSH) methods. Enhancements include head-wise attention amplification (HeadAmp), diversity-promoting attention (DivAttn), and cross-level attention via Nyström approximation (CrossLevel). Percentages in parentheses show improvement ($+$) or decline (--) relative to base GAT.
Method $AC$ $F1$ $E$ $R$
@@
We compare different GNN architectures within GAIS (specifically the DM method) to thoroughly assess our method’s design, where only the GNN component is replaced while maintaining the same graph construction and IS mechanisms, as shown in Table 7. Traditional GNN architectures (GCN, GraphSAGE, GIN (Kipf and Welling, 2017; Hamilton et al., 2018; Xu et al., 2019)) achieve comparable effectiveness scores ($0.793-0.795$), while attention-based approaches like GATv2 show similar performance ($0.795$). The Graph Transformer achieves the highest effectiveness among the baseline GNN architectures ($0.796$). However, GAIS with our enhanced GAT architecture consistently outperforms all alternatives across all considered metrics ($AC=0.837;F1=0.680;E=0.807;R=0.963$). This supports the importance of our added enhancements which help to better identify informative instances compared to base GNN architectures. Although GNNs can provide viable node importance scores through learned representations, considering attention mechanisms explicitly provides more useful signals for selection.
@@
Bold: highest; italics: 2nd highest; underline: 3rd highest value in each row distance-based mini-batching (DM) GAIS RIS 1: Cavalcanti and Soares (2020); BDIS: Chen et al. (2022b); CNN—EGDIS: Malhat et al. (2020); EIS-AS: Gong et al. (2021); CIS-BE: Moran et al. (2022); S2EJR-2: Sánchez-Fernández et al. (2024)

### v3

@@ 1 Introduction
Traditional IS methods rely mostly on nearest neighbor computations, falling into three groups: condensation (keeping instances near decision boundaries), edition (removing noisy instances), and hybrid approaches (Cano et al., 2003; García et al., 2012; Saha et al., 2022). While these work well for moderately sized datasets, they struggle to capture complex relationships in high-dimensional spaces. Their need for pairwise distance calculations results in quadratic complexity that becomes prohibitive for large-scale datasets, often taking several hours of preprocessing time on datasets with millions of instances.
Graph-based methods offer a promising alternative by representing data relationships through node connections (Goyal and Ferrara, 2018). These approaches can effectively capture structural information through proximity graphs and neighborhood networks, revealing dataset patterns that distance-based methods might miss. Recent advances in graph sampling and reduction (Chen et al., 2022) show their potential to handle large-scale computational challenges while retaining key data characteristics.
@@
Graph Attention Networks (GATs) (Veličković et al., 2018) present an unexplored direction for IS through their ability to learn importance weights for node neighborhoods dynamically. The attention mechanisms in GATs can compute instance importance based on local structural properties, providing an interpretable approach to identify informative data points. However, two questions remain: (1) How can we efficiently construct meaningful graph representations from large tabular datasets? (2) How can we utilize attention mechanisms to identify truly informative instances while keeping computational costs feasible?
@@ 2.1 Instance Selection
Traditional IS methods primarily rely on nearest neighbor computations to evaluate instance relationships and importance, falling into three main categories: condensation (retaining instances near decision boundaries), edition (removing noisy or borderline instances), and hybrid approaches (García et al., 2012). While various improvements have been proposed, including density-based and scalability-focused approaches (Carbonera and Abel, 2018; García-Osorio et al., 2010), recent work has introduced several innovative methodologies that enhance or reformulate how instance relationships are evaluated.
Density-based approaches have evolved from local to global analysis, as demonstrated by GDIS and EGDIS (Malhat et al., 2020), which evaluate instance importance through global density patterns and neighbor relationships. The ranking-based RIS algorithm (Cavalcanti and Soares, 2020) introduces a scoring mechanism that considers relationships between instances, using normalized exponential transform of distances weighted by class membership. More theoretically grounded approaches such as BDIS (Chen et al., 2022) employ Bayesian decision theory to categorize instances as reducible, irreducible, or deleterious, while leveraging percolation theory to identify local homogeneous clusters—groups of instances with the same label connected by nearest neighbors, where the cluster size reflects the probability of class consistency. Alternative frameworks like CIS (Moran et al., 2022) employ reinforcement learning and clustering, where an agent learns to select valuable instance clusters through Q-learning and curiosity-driven exploration. Recent work has explored novel theoretical frameworks, including approval-based multi-winner voting (Sánchez-Fernández et al., 2024), which models instances as both voters and candidates in an election system, and evidence theory-based approaches like EIS (Gong et al., 2021), which evaluates instance importance through mass functions and Dempster’s rule of combination.
@@ 2.2 Graph-based Selection
Graph reduction techniques (Chen et al., 2022), while successful in general graph summarization tasks, have seen limited application in IS. As shown in Figure 2, these methods merge nodes or modify edge structures. When applied to IS, these approaches typically require mapping reduced nodes back to original instances, often by selecting representatives from merged clusters. This indirect correspondence makes these methods less suitable for IS tasks compared to graph sampling approaches that maintain direct relationships to the original instances.
@@
The effectiveness of graph-based methods for IS has been demonstrated by Rustamov et al. (2025), who evaluated multiple graph-based approaches across diverse classification datasets, showing their potential to achieve substantial data reduction while maintaining model performance. In our previous work (Rustamov et al., 2024), we explored using GATs for IS. While this approach showed promise in identifying important instances through attention mechanisms, it was limited by its reliance on confidence scores as importance measures and faced significant scalability issues with large datasets due to its inefficient chunking-based graph construction that still required substantial memory overhead. In this paper, we address these limitations by introducing efficient graph construction methods—distance-based mini-batching and LSH approaches—that significantly reduce computational complexity, alongside a dual-perspective importance scoring mechanism that leverages attention weights to more effectively identify informative instances.
@@ 2.3 Scaling Challenges in Graph-based Methods
While graph-based methods show promise for IS, they face significant scaling challenges as datasets grow in size and dimensionality. The primary bottleneck lies in the graph construction phase, particularly for k-nearest neighbor (KNN) graphs that form the foundation of many graph-based approaches. Constructing KNN graphs requires computing pairwise distances between instances, resulting in $O(n^{2})$ computational complexity and substantial memory requirements for storing the resulting adjacency matrices (Li et al., 2025).
@@ 3.1.1 Distance-based Mini-Batching Method
The sampling process maintains a set $\mathcal{S}_{k}$ of sampled instances, updated after each mini-batch: $\mathcal{S}_{k}=\mathcal{S}_{k-1}\cup B_{k}$ with $\quad\mathcal{S}_{0}=\emptyset.$ Subsequent mini-batches are sampled from $\mathcal{D}_{k}=\mathcal{D}\setminus\mathcal{S}_{k-1}$. Let $C_{c}=\{(x_{i},y_{i})\in\mathcal{D}:y_{i}=c\}$ denote the set of all instances from class $c$. To ensure each class is represented in every mini-batch, we enforce $\min_{c}\,\left|\,B_{k}\cap C_{c}\right|\;\geq\;1.$
@@ 3.1.2 Locality-Sensitive Hashing-based Method
Figure 5: High-level overview of the Multi-Level LSH (ML-LSH) approach. ML-LSH processes instances through multiple hierarchical levels (Level 11 to $M$), where each level follows a four-step process: (1) Random projections using hash functions, (2) Hash code generation (e.g., $h(x)=\text{sign}(P^{T}x)$), (3) Bucket assignment grouping similar instances, and (4) Edge connection within buckets. Lower levels (Level 11) capture coarse relationships using fewer hash tables and projections, while higher levels (Level $M$) employ exponentially increasing hashing parameters to refine local similarities. The final graph combines edge sets from all levels, with each level treated independently to capture different relationship granularities.
@@
$\mathrm{Ham}\left(h_{1},\;h_{2}\right)\;\leq\;1,\quad\lvert B_{h_{1}}\rvert+\lvert B_{h_{2}}\rvert\;<\;\gamma.$
@@
$L_{v,m}\;=\;L_{v}\cdot 2^{m},\quad k_{v,m}\;=\;k_{v}\cdot 2^{m},\quad\theta_{v,m}\;=\;\frac{\theta_{v}}{\,2^{m}}.$
@@ 3.2 Graph Attention Network (GAT)
GATs provide an ideal foundation for IS due to their ability to learn adaptive node relationships through attention mechanisms. While traditional GATs excel at capturing local structural patterns, IS requires a more nuanced understanding of both local and global data relationships. We build upon the foundational GAT architecture (Veličković et al., 2018) by introducing four key enhancements: (1) head-wise attention amplification through learnable scaling factors, (2) diversity-promoting attention that encourages distinct patterns across attention heads, (3) multi-level/multi-view attention mechanisms that process our graph representations at multiple scales, and (4) cross-level attention via Nyström approximation for efficient global information exchange.
@@ 3.3.2 Selection Strategies
Table 1: Dataset statistics with domain, number of instances ($N$), features ($D$), and classes ($K$). Dataset size indicators: small ($N<10,000$), medium ($10,000\leq N\leq 100,000$), large ($N>100,000$).
@@ 4.1 Datasets and Models
This study employs diverse classification datasets from the UCI Machine Learning Repository [^1] UCI Machine Learning Repository. Kelly, M., Longjohn, R., Nottingham, K. (2023). Available at: https://archive.ics.uci.edu (accessed 4 October 2024). to evaluate our proposed GAIS method. Table 1 summarizes these datasets, including their domain, number of instances ($N$), features ($D$), and classes ($K$). The datasets vary widely in size (from 303 to five million instances), dimensionality (two to 784 features), and class distributions (balanced and imbalanced), spanning domains such as healthcare, game theory, and bioinformatics. We evaluate our method using Random Forest and Logistic Regression as representative non-linear and linear classifiers, respectively.
@@ 4.5 Evaluation Metrics
Accuracy ACAC = (TP + TN) / (TP + TN + FP + FN)
Precision PRPR = TP / (TP + FP)
Recall RERE = TP / (TP + FN)
@@ 5 Results and Discussions
Method ACAC PRPR RERE $F1$ $E$ $E_{F1}$ $R$
Small Datasets (<10,000<10,000)
@@ 5.1 Comparing GAIS Variants
For large datasets, the DM method demonstrates the highest efficiency with the lowest construction and training times (GCT = 2.55s, GTT = 10.31s), making it particularly suitable for large-scale applications where computational resources are limited. The more complex MVML-LSH and CMVML-LSH methods show significantly higher computational requirements (GCT >> 35s, GTT >> 80s), though this should be weighed against their enhanced performance metrics. Importantly, the original ML training time (70.64s) is substantially higher than the training time on selected instances (0.28-0.82s), highlighting the significant computational savings achieved through IS, regardless of the chosen method.
@@ 5.3 Component Analysis
Figure 8 illustrates the relationship between mini-batch configuration and test effectiveness in the DM method for large datasets. We test configurations up to $w_{\max}$ instances per mini-batch to identify natural performance plateaus. The results demonstrate that both the number of mini-batches and instances per batch have a positive impact on effectiveness, though this improvement shows diminishing returns. As the number of mini-batches increases from 1 to 9, the effectiveness consistently improves across all instance sizes, with the most substantial gains observed in the transition from 1 to 5 mini-batches. Similarly, larger batch sizes generally yield better effectiveness, though the improvement plateaus around $7000$ instances, particularly with higher numbers of mini-batches. This plateau effect is most noticeable when using 7-9 mini-batches, suggesting that beyond these configurations, additional computational resources might not yield proportional performance benefits. These findings provide practical guidance for implementing the DM method, indicating that moderate configurations (5-7 mini-batches with 500070005000-7000 instances per batch) might offer an optimal balance between computational efficiency and effectiveness. The empirical analysis demonstrates that effectiveness plateaus around $7000$ instances per mini-batch, providing direct guidance for $w_{\max}$ selection. This plateau effect validates that computational bounds can be set without sacrificing performance.
Figure 8: Impact of mini-batch configuration on test effectiveness in the distance-based mini-batch (DM) method for large datasets (>>100K instances). Results show diminishing returns beyond 7,000 instances per mini-batch and 5-7 mini-batches, providing empirical guidance for setting the computational constraint $w_{\max}$ while maintaining performance.
@@
Figure 9: Effectiveness versus sampling percentage by dataset size categories: small (<<10K instances), medium (10K–100K instances), and large (>>100K instances). Performance plateaus occur at 5%, 10%, and 40% sampling for large, medium, and small datasets respectively, with the ”Diminishing Returns Zone” indicating where additional sampling provides minimal benefit.
@@
Method Sep $\uparrow$ Pur $\uparrow$ $R@5$ $\uparrow$ $\rho$ $\uparrow$ $\bar{B}$ $\downarrow$
Small Datasets (<10,000<10,000)
@@
Figure 11: Multi-level and multi-view analysis showing relative improvement (in %) in test effectiveness for LSH-based graph construction variants. (a) Multi-level analysis: Improvement from using hierarchical levels compared to single-level baseline, across small (<<10K instances), medium (10K–100K instances), and large (>>100K instances) datasets. (b) Multi-view analysis: Improvement from using multiple views compared to single-view baseline. Solid lines show average improvements across all datasets in each size category, while dashed lines represent improvements for datasets performing below median effectiveness within each category to highlight challenging datasets.
Figure 12 reveals how different IS strategies impact test effectiveness across varying retention rates. The balanced class-aware strategy demonstrates better performance at higher retention rates ($\rho=0.9$$0.3$), likely due to its ability to maintain equal class representation, which is particularly beneficial when selecting a larger subset of instances. However, this advantage diminishes as the retention rate decreases, with both the global top-percentage and proportional class-aware approaches outperforming the balanced strategy for $\rho<0.2$. All three strategies exhibit peak performance around the $\rho=0.06$$0.04$ range, after which effectiveness begins to decline, most notably in the top-percentage and balanced approaches, while the proportional strategy shows slightly more resilience at lower retention rates. This suggests that while class-aware selection mechanisms are crucial for maintaining model performance when selecting more instances, preserving the original class distribution (proportional strategy) becomes increasingly important as the retention rate decreases toward its optimal level.
Figure 12: Test effectiveness versus retention rate $\rho$ for three GAIS selection strategies: top-% (global selection), class-aware balanced (equal per class), and class-aware proportional (maintains class distribution). Inset shows optimal range ($\rho=0.1$$0.01$) where proportional strategy performs best, while balanced strategy excels at higher retention rates.
@@
Table 6: Ablation study of GAT enhancements for distance-based mini-batching (DM) and multi-level locality-sensitive hashing (ML-LSH) methods. Enhancements include head-wise attention amplification (HeadAmp), diversity-promoting attention (DivAttn), and cross-level attention via Nyström approximation (CrossLevel). Percentages in parentheses show improvement (++) or decline (-) relative to base GAT.
Method ACAC $F1$ $E$ $R$
@@
We compare different GNN architectures within GAIS (specifically the DM method) to thoroughly assess our method’s design, where only the GNN component is replaced while maintaining the same graph construction and IS mechanisms, as shown in Table 7. Traditional GNN architectures (GCN, GraphSAGE, GIN (Kipf and Welling, 2017; Hamilton et al., 2017; Xu et al., 2019)) achieve comparable effectiveness scores (0.7930.7950.793-0.795), while attention-based approaches like GATv2 (Brody et al., 2022) show similar performance ($0.795$). The Graph Transformer (Shi et al., 2021) achieves the highest effectiveness among the baseline GNN architectures ($0.796$). However, GAIS with our enhanced GAT architecture consistently outperforms all alternatives across all considered metrics ($AC=0.837;F1=0.680;E=0.807;R=0.963$). This supports the importance of our added enhancements which help to better identify informative instances compared to base GNN architectures. Although GNNs can provide viable node importance scores through learned representations, considering attention mechanisms explicitly provides more useful signals for selection.
@@
Bold: highest; italics: 2nd highest; underline: 3rd highest value in each row distance-based mini-batching (DM) GAIS RIS 1: Cavalcanti and Soares (2020); BDIS: Chen et al. (2022); CNN—EGDIS: Malhat et al. (2020); EIS-AS: Gong et al. (2021); CIS-BE: Moran et al. (2022); S2EJR-2: Sánchez-Fernández et al. (2024)